OpenStudy (anonymous):

log2 5 +log2 a=log2 17??

OpenStudy (watergirl):



OpenStudy (anonymous):

what is this @WATERGIRL

OpenStudy (watergirl):



OpenStudy (whpalmer4):

\[\log_2 5 + \log_2 a = \log_2 17\]Use log property that \(\log a + \log b = \log a*b\)\[\log_2 5a = \log_2 17\]Solve for \(a\)